Annapolis Company has two service departments (Computer Operations & Maintenance Services). Annapolis has two production departments (Mixing Department & Packaging Department.) Annapolis uses a direct allocation method where service departments are allocated only to production departments. All allocations are based on total employees. Computer Operations has costs of $105,000 and Maintenance Services has costs of $160,000 before any allocations. What amount of Maintenance Services total cost is allocated to the Packaging Department? (round to closest whole dollar) Employees are: Computer Operations 2 Maintenance Services 2 Mixing Department 7 Packaging Department 6
Answer #1

Total employees for production department = 7+6 = 13

Maintenance cost allocated to packaging department = Maintenance cost/Total employee*employee for packaging department

= 160000*6/13

Maintenance cost allocated to packaging department = 73846